# Artificial Intelligence in Image-based Cardiovascular Disease Analysis: A Comprehensive Survey and Future Outlook

### v3

@@ 3.2 Non-vessel Structures
3.2.1 Atria
@@
Atrial Diseases. We examine several common diseases related to the atria [126, 127]. A key condition, Atrial fibrillation (AF), is a prevalent cardiac arrhythmia marked by irregular and rapid atrial electrical activity [128, 129]. Instead of regular contractions, the atria fibrillate or quiver, resulting in an irregular heartbeat, leading to symptoms like palpitations, shortness of breath, fatigue, and dizziness [130]. Crucially, AF heightens the risk of stroke due to potential blood clot formation in the atria [131].
For detecting AF-related abnormalities, primary imaging techniques include ECG and cardiac MRI. ECG is preferred for its real-time imaging, while cardiac MRI, especially Late Gadolinium Enhancement (LGE) MRI, offers high-resolution images vital for assessing structural changes and fibrosis. Deep learning algorithms trained on extensive ECG datasets can effectively identify AF patterns [131], achieving high accuracy that often matches or exceeds human experts [132]. As shown in Figure 1 (a), the AI analysis pipeline for AF ablation involves tasks like LA cavity and wall segmentation, scar segmentation, quantification, and applications such as locating ablation gaps from LGE MRI [133].
Quantifying atrial abnormalities often involves measuring atrial volume, detecting fibrotic tissue, and assessing atrial wall thickness. These features are derived using various image processing methods and quantified through shape and intensity metrics. While these features are generally specific and sensitive, particularly with advanced imaging, the variability in AF manifestations challenges standardization. For comprehensive information on AI methods in AF research, including LA cavity, wall, and scar segmentation from LGE MRI, refer to a recent survey [20].
@@
Atrial Enlargement (AE), another prevalent atrial condition, involves the enlargement of one or both atria, often resulting from chronic issues like hypertension and heart valve diseases [134]. AE is linked to an increased risk of arrhythmias and heart failure [135]. Various imaging modalities, including ECG, echocardiography, cardiac MRI, and CT scans, are employed to detect and assess AE. ECG stands out for its real-time imaging and widespread availability, effectively identifying atrial size and function. Doppler echocardiography complements ECG by providing detailed hemodynamic data. Cardiac MRI and CT scans are valued for their high spatial resolution, enabling precise quantification of atrial volume with heightened sensitivity and specificity. Deep learning technologies, particularly Convolutional Neural Networks (CNNs), have been trained on extensive datasets to recognize ECG patterns indicative of AE [135]. These models offer clinicians a rapid and accurate preliminary diagnosis based on ECG data. The quantification of AE involves measuring atrial dimensions, volume, and area, as well as assessing shape descriptors and functional parameters like atrial emptying fractions. While MRI and CT provide high specificity and sensitivity for these measurements, echocardiography may encounter challenges related to image quality and variability among operators [136]. The integration of these imaging techniques with AI advancements offers a powerful toolkit for the early detection and quantification of atrial enlargement.
3.2.2 Ventricle
@@
The heart ventricles, comprising the lower two chambers, are vital components of the heart’s anatomy, as illustrated in Figure 1 (b). The left ventricle, recognized as the heart’s largest and strongest chamber, plays a critical role in receiving oxygen-rich blood from the left atrium and pumping it into the systemic circulation. This function is essential for supplying oxygen and nutrients to the body’s organs and tissues. On the other hand, the right ventricle is responsible for receiving oxygen-depleted blood from the right atrium and directing it into the pulmonary circulation [137, 138]. In this section, we explore various diseases related to the ventricles and delve into how AI technologies are being utilized to analyze these conditions [139]. This discussion will highlight the intersection of advanced technological applications and cardiology, demonstrating the impact of AI in the understanding and treatment of ventricular diseases.
Ventricular Diseases. We delve into various anomalies related to the structure and function of the ventricles, each presenting unique characteristics and clinical implications [140]. For instance, abnormal heart rhythms can originate from the ventricles, ranging from benign conditions like premature ventricular contractions (PVCs) to more serious disorders such as ventricular tachycardia or fibrillation [141]. These arrhythmias can compromise the heart’s normal pumping function, potentially leading to life-threatening situations [142]. In contrast, Ventricular Septal Defect (VSD) is a congenital heart defect characterized by an abnormal opening in the septum dividing the ventricles. This defect allows the mixing of oxygen-rich and oxygen-poor blood, impacting the heart’s efficiency. The size and severity of VSDs can vary [12].
@@
As indicated in Figure 1 (b), the Ejection Fraction (EF) is a key measure of the heart’s pumping efficiency, particularly of the left ventricle. It is critical for identifying patients at risk of heart dysfunctions such as heart failure [143]. EF represents the proportion of blood ejected from the left ventricle with each heartbeat [144]. Normally, the left ventricle effectively pumps out over half of its blood content with each beat. A reduced EF signifies impaired heart pumping function [145]. A notable recent study introduced a model using Graph Neural Networks (GNNs) to estimate EF from echocardiography videos. This model, EchoGNN, demonstrates EF prediction accuracy comparable to current leading methods and offers crucial explainability, addressing the high variability often seen in observer interpretations [146].
3.2.3 Myocardium
@@
The myocardium, the heart’s muscular tissue and its middle layer, is crucial for the organ’s function, as illustrated in Figure 1 (c). Comprising specialized cardiac muscle cells called cardiomyocytes, the myocardium is responsible for the heart’s contractions and relaxations, essential for pumping blood throughout the body. Its role in maintaining heart function and ensuring proper blood circulation is fundamental [147]. The myocardium receives its supply of oxygenated blood from the coronary arteries, which nourish the myocardial cells with oxygen and essential nutrients. With each heartbeat, the myocardium contracts, effectively propelling blood from the heart chambers into the circulatory system. Various diseases can impair the myocardium, affecting its ability to function effectively. In the following discussion, we explore several common myocardial conditions and their implications [148]. These diseases illustrate the myocardium’s vulnerability and the importance of maintaining its health for overall cardiovascular well-being.
@@
Myocardial Diseases. Myocardial Infarction (Heart Attack) is a critical myocardial function-related disease, arising when blood flow to a part of the myocardium is obstructed, typically by a clot in the coronary arteries [149, 150]. This blockage can cause damage or death to the affected myocardial area, leading to chest pain, shortness of breath, and potentially severe complications [151]. Innovations in deep learning have enabled the use of non-enhanced cardiac MRI to detect and quantify chronic myocardial infarction, potentially reducing reliance on gadolinium contrast injections [152]. Additionally, [153] developed machine learning models that combine cardiac troponin levels with clinical features to evaluate an individual’s myocardial infarction risk. Myocardial perfusion MRI, a noninvasive imaging technique, plays a vital role in detecting ischemic heart disease with high accuracy. Figure 1 (c) demonstrates how AI models provide detailed heart blood flow images, crucial for diagnosing conditions associated with impaired blood supply to the heart muscle, thereby aiding in effective cardiac health management.
Cardiomyopathy, another prevalent structural myocardial anomaly, encompasses diseases that affect the heart muscle, potentially impacting ejection fraction [154]. A notable subtype, Hypertrophic Cardiomyopathy, is a genetic disorder marked by abnormal myocardial thickening, especially in the left ventricle [155]. This thickening can hinder effective blood pumping, manifesting as chest pain, shortness of breath, fainting, and arrhythmias [156]. Hypertrophic cardiomyopathy is especially significant as a leading cause of sudden cardiac arrest in young people.
3.3 Vessel Structure and Diseases
@@
3.3.1 Aorta
@@
The aorta, the human body’s largest artery, is a crucial component of the circulatory system, as depicted in Figure 1 (d). Arising from the heart’s left ventricle, it distributes oxygenated blood throughout the body, providing essential nutrients and oxygen to various organs and tissues [157]. Structurally, the aorta is segmented into the ascending aorta, the aortic arch, and the descending aorta, each playing a distinct role in blood circulation [158]. This section offers an overview of some common diseases associated with the aorta. Understanding these conditions is essential for comprehending the aorta’s critical role in maintaining overall cardiovascular health and identifying potential risks and complications that may arise from aortic disorders.
@@
Aortic Diseases. Aortic Stenosis, as outlined in [159], is a condition marked by the narrowing of the aortic valve, impeding blood flow from the left ventricle to the aorta. Common in adults due to valve calcification or age-related degeneration, it can cause chest pain, shortness of breath, fatigue, and fainting [160]. The evaluation of heart function is critically enhanced through the analysis of hemodynamic flow parameters [161, 162]. Four-dimensional (4D) flow MRI, a significant advancement in cardiovascular imaging, is extensively discussed in [163]. This technology enables the detailed study of flow dynamics in vivo, essential for understanding and quantifying the parameters relevant to aortic disease [164]. The importance of discerning these flow dynamics lies in their potential to contribute to or worsen vascular diseases, as detailed in [165]. Notably, the innovative 4DFlowNet, introduced in [166], demonstrates the capability to enhance the spatial resolution of 4D flow MRI, showing its effectiveness in both phantom and actual patient data.
Aortic Dissection, as detailed in [167], is a critical condition characterized by a tear in the aorta’s inner wall. This tear permits blood flow between the layers of the aortic wall, potentially causing separation of these layers and the formation of a false lumen. Aortic dissection typically presents with abrupt and severe chest or back pain and is a medical emergency requiring immediate attention. Its life-threatening nature demands prompt treatment. For further insights into AI’s role in aortic dissection imaging, refer to the comprehensive review [168]. This review sheds light on three key aspects: the processes involved in screening for aortic dissection, the application of image segmentation techniques, and the use of predictive analytics in determining patient outcomes. The review encapsulates the significant advancements AI has made in the field, providing a thorough understanding of its applications in diagnosing and managing aortic dissection.
3.3.2 Coronary Arteries
@@
Coronary arteries, as illustrated in Figure 1 (e), play a crucial role in the human circulatory system by delivering oxygen-rich blood to the heart muscle [169, 170, 171, 172, 173]. These arteries branch out from the aorta—the body’s main artery—and envelop the heart, forming an extensive network of vessels essential for cardiac function. Diseases affecting the coronary arteries can lead to myocardial ischemia, a condition characterized by reduced blood flow to the heart muscle [174, 175, 176]. Chronic ischemia or its associated complications can progressively impair the heart’s pumping efficiency, posing significant health risks [177]. Early detection and timely medical intervention are crucial for conditions related to coronary arteries, as they can be life-threatening if left unaddressed [178].
In the subsequent discussion, we delve into the realm of AI-driven analysis of diseases associated with coronary arteries. This exploration includes a comprehensive overview of the latest AI methodologies applied in the detection, diagnosis, and management of coronary artery diseases.
Figure 5: Examples of connections of non-vessel and vessel structures. Top: A relation example of the Coronary Artery with corresponding LV regions for providing blood. (LAD: left anterior descending; RCA: right CA; LCX: left circumflex) [3, 179] and Visualization of Myocardial Perfusion from Coronary Anatomy[180]. Bottom (left): Determination of Myocardial Perfusion Territory [181]. Bottom (right): PET/CT assessing perfusion and Coronary Artery anatomy [182].
Coronary Artery Diseases (CAD). CAD, the most prevalent form of heart disease, arises when coronary arteries are narrowed or blocked due to fatty plaque buildup [183]. This condition can significantly reduce or completely obstruct blood flow to the heart, leading to angina, manifested as chest pain or discomfort. CAD may also precipitate a heart attack, or myocardial infarction, where a part of the heart muscle sustains damage or death due to insufficient blood supply [184].
Advanced stages of CAD often necessitate invasive treatments like Percutaneous Coronary Intervention (PCI) or coronary angioplasty, which involve widening narrowed arteries using a balloon device and potentially placing a stent to keep the artery open. In more severe cases with multiple blockages, Coronary Artery Bypass Grafting (CABG) surgery might be performed, creating alternative pathways for blood flow using vessels from other parts of the body [185]. Timely and effective management of CAD is crucial. Below, we discuss significant research challenges associated with CAD.
Over the years, AI techniques have become extensively utilized in the field of Coronary Artery Disease detection [186]. For example, Coronary Artery Stenosis is a predominant structural concern in coronary arteries. The detection, characterization, and monitoring of stenosis severity and plaque development are vital, as these conditions can lead to serious events like heart attacks [83]. Medical imaging is central to these tasks [187, 170, 188, 189, 171, 190, 191, 192]. Notably, interpretations of coronary computed tomography angiography (CTA) by less experienced clinicians often overestimate the severity of stenosis compared to expert analyses [50]. AI solutions present a promising alternative. A recent study [193] assessed AI-augmented quantitative coronary computed tomography (AI-QCT) angiography for detecting and grading coronary stenoses.

### v4

@@ 3.3 Non-vessel Structures
In this section, we provide a summary of CVD analysis focusing on the Non-vessel structures.
3.3.1 Atria and Atrial Diseases
In this section, we introduce the main structure and function of the atria, as well as AI technologies used for analyzing related diseases.
3.3.1.1 Atria
@@
3.3.1.2 Atrial Diseases
@@
We examine several common diseases related to the atria [wang2022deep, ter2023juvenile]. A key condition, Atrial fibrillation (AF), is a prevalent cardiac arrhythmia marked by irregular and rapid atrial electrical activity [raicea2021giant, wang2022current]. Instead of regular contractions, the atria fibrillate or quiver, resulting in an irregular heartbeat, leading to symptoms like palpitations, shortness of breath, fatigue, and dizziness [makynen2022wearable]. Crucially, AF heightens the risk of stroke due to potential blood clot formation in the atria [raghunath2021deep]. For detecting AF-related abnormalities, primary imaging techniques include ECG and cardiac MRI. ECG is preferred for its real-time imaging, while cardiac MRI, especially Late Gadolinium Enhancement (LGE) MRI, offers high-resolution images vital for assessing structural changes and fibrosis. Deep learning algorithms trained on extensive ECG datasets can effectively identify AF patterns [raghunath2021deep], achieving high accuracy that often matches or exceeds human experts [murat2021review]. As shown in Figure 1 (a), the AI analysis pipeline for AF ablation involves tasks like LA cavity and wall segmentation, scar segmentation, quantification, and applications such as locating ablation gaps from LGE MRI [li2020atrial]. Atrial Enlargement (AE), another prevalent atrial condition, involves the enlargement of one or both atria, often resulting from chronic issues like hypertension and heart valve diseases [hsu2022machine]. AE is linked to an increased risk of arrhythmias and heart failure [jiang2020detection]. Various imaging modalities, including ECG, echocardiography, cardiac MRI, and CT scans, are employed to detect and assess AE. ECG stands out for its real-time imaging and widespread availability, effectively identifying atrial size and function. Doppler echocardiography complements ECG by providing detailed hemodynamic data. Cardiac MRI and CT scans are valued for their high spatial resolution, enabling precise quantification of atrial volume with heightened sensitivity and specificity. Deep learning technologies, particularly Convolutional Neural Networks (CNNs), have been trained on extensive datasets to recognize ECG patterns indicative of AE [jiang2020detection]. These models offer clinicians a rapid and accurate preliminary diagnosis based on ECG data. While MRI and CT provide high specificity and sensitivity for these measurements, echocardiography may encounter challenges related to image quality and variability among operators [kuchynka2015role]. The integration of these imaging techniques with AI advancements offers a powerful toolkit for the early detection and quantification of atrial enlargement.
3.3.2 Ventricle and Ventricular Diseases
In this section, we introduce the main structure and function of the ventricle, as well as AI technologies used for analyzing related diseases.
3.3.2.1 Ventricle
@@
The heart ventricles, comprising the lower two chambers, are vital components of the heart’s anatomy, as illustrated in Figure 1 (b). The left ventricle, recognized as the heart’s largest and strongest chamber, plays a critical role in receiving oxygen-rich blood from the left atrium and pumping it into the systemic circulation. This function is essential for supplying oxygen and nutrients to the body’s organs and tissues. On the other hand, the right ventricle is responsible for receiving oxygen-depleted blood from the right atrium and directing it into the pulmonary circulation [pantelidis2023deep, zhao2022deep].
3.3.2.2 Ventricular Diseases
@@
We delve into various anomalies related to the structure and function of the ventricles, each presenting unique characteristics and clinical implications [olaisen2023automatic]. For instance, abnormal heart rhythms can originate from the ventricles, ranging from benign conditions like premature ventricular contractions (PVCs) to more serious disorders such as ventricular tachycardia or fibrillation [madan2022hybrid]. These arrhythmias can compromise the heart’s normal pumping function, potentially leading to life-threatening situations [kolk2023machine]. In contrast, Ventricular Septal Defect (VSD) is a congenital heart defect characterized by an abnormal opening in the septum dividing the ventricles. This defect allows the mixing of oxygen-rich and oxygen-poor blood, impacting the heart’s efficiency. The size and severity of VSDs can vary [yang2023application]. As indicated in Figure 1 (b), the Ejection Fraction (EF) is a key measure of the heart’s pumping efficiency, particularly of the left ventricle. It is critical for identifying patients at risk of heart dysfunctions such as heart failure [kusunose2021standardize]. EF represents the proportion of blood ejected from the left ventricle with each heartbeat [liu2021deep]. Normally, the left ventricle effectively pumps out over half of its blood content with each beat. A reduced EF signifies impaired heart pumping function [akerman2023automated]. A notable recent study introduced a model using Graph Neural Networks (GNNs) to estimate EF from echocardiography videos. This model, EchoGNN, demonstrates EF prediction accuracy comparable to current leading methods and offers crucial explainability, addressing the high variability often seen in observer interpretations [mokhtari2022echognn].
3.3.3 Myocardium and Myocardial Diseases
In this section, we introduce the main structure and function of the myocardium, as well as AI technologies used for analyzing related diseases.
3.3.3.1 Myocardium
@@
The myocardium, the heart’s muscular tissue and its middle layer, is crucial for the organ’s function, as illustrated in Figure 1 (c). Comprising specialized cardiac muscle cells called cardiomyocytes, the myocardium is responsible for the heart’s contractions and relaxations, essential for pumping blood throughout the body. Its role in maintaining heart function and ensuring proper blood circulation is fundamental [rinaldi2022invasive]. The myocardium receives its supply of oxygenated blood from the coronary arteries, which nourish the myocardial cells with oxygen and essential nutrients. With each heartbeat, the myocardium contracts, effectively propelling blood from the heart chambers into the circulatory system. Various diseases can impair the myocardium, affecting its ability to function effectively. In the following discussion, we explore several common myocardial conditions and their implications. These diseases illustrate the myocardium’s vulnerability and the importance of maintaining its health for overall cardiovascular well-being.
3.3.3.2 Myocardial Diseases
@@
We examine several common diseases related to the Myocardium. Myocardial Infarction (Heart Attack) is a critical myocardial function-related disease, arising when blood flow to a part of the myocardium is obstructed, typically by a clot in the coronary arteries [zheng2019explainable, ibanez20182017]. This blockage can cause damage or death to the affected myocardial area, leading to chest pain, shortness of breath, and potentially severe complications [cho2020artificial]. Innovations in deep learning have enabled the use of non-enhanced cardiac MRI to detect and quantify chronic myocardial infarction, potentially reducing reliance on gadolinium contrast injections [zhang2019deep]. Additionally, [doudesis2023machine] developed machine learning models that combine cardiac troponin levels with clinical features to evaluate an individual’s myocardial infarction risk. Myocardial perfusion MRI, a noninvasive imaging technique, plays a vital role in detecting ischemic heart disease with high accuracy. Figure 1 (c) demonstrates how AI models provide detailed heart blood flow images, crucial for diagnosing conditions associated with impaired blood supply to the heart muscle, thereby aiding in effective cardiac health management. Cardiomyopathy, another prevalent structural myocardial anomaly, encompasses diseases that affect the heart muscle, potentially impacting ejection fraction [vukadinovic2023deep]. A notable subtype, Hypertrophic Cardiomyopathy, is a genetic disorder marked by abnormal myocardial thickening, especially in the left ventricle [guo2022artificial]. This thickening can hinder effective blood pumping, manifesting as chest pain, shortness of breath, fainting, and arrhythmias [wang2023deephcd]. Hypertrophic cardiomyopathy is especially significant as a leading cause of sudden cardiac arrest in young people [doi:10.1016/j.jacep.2018.11.004].
3.4 Vessel Structures
@@
3.4.1 Aorta and Aortic Diseases
In this section, we introduce the main structure and function of the aorta, as well as AI technologies used for analyzing related diseases.
3.4.1.1 Aorta
@@
The aorta, the human body’s largest artery, is a crucial component of the circulatory system, as depicted in Figure 1 (d). Arising from the heart’s left ventricle, it distributes oxygenated blood throughout the body, providing essential nutrients and oxygen to various organs and tissues [chandrashekar2021deep]. Structurally, the aorta is segmented into the ascending aorta, the aortic arch, and the descending aorta, each playing a distinct role in blood circulation [stonko2023review]. This section offers an overview of some common diseases associated with the aorta. Understanding these conditions is essential for comprehending the aorta’s critical role in maintaining overall cardiovascular health and identifying potential risks and complications that may arise from aortic disorders.
3.4.1.2 Aortic Diseases
@@
We examine several common diseases related to the Aorta. Aortic Stenosis, as outlined in [holste2023severe], is a condition marked by the narrowing of the aortic valve, impeding blood flow from the left ventricle to the aorta. Common in adults due to valve calcification or age-related degeneration, it can cause chest pain, shortness of breath, fatigue, and fainting [elvas2023ai]. The evaluation of heart function is critically enhanced through the analysis of hemodynamic flow parameters [ramaekers2023clinician, zhuang2021role]. Four-dimensional (4D) flow MRI, a significant advancement in cardiovascular imaging, is extensively discussed in [bissell20234d]. This technology enables the detailed study of flow dynamics in vivo, essential for understanding and quantifying the parameters relevant to aortic disease [juffermans20224d]. The importance of discerning these flow dynamics lies in their potential to contribute to or worsen vascular diseases, as detailed in [catapano20204d]. Notably, the innovative 4DFlowNet, introduced in [ferdian20204dflownet], demonstrates the capability to enhance the spatial resolution of 4D flow MRI, showing its effectiveness in actual patient data.
3.4.2 Coronary Arteries and Coronary Artery Diseases
In this section, we introduce the main structure and function of the coronary arteries, as well as AI technologies used for analyzing related diseases.
3.4.2.1 Coronary Arteries
@@
Coronary arteries, as illustrated in Figure 1 (e), play a crucial role in the human circulatory system by delivering oxygen-rich blood to the heart muscle [awad2014public, li2001magnetic, pennell1993magnetic, de2023framing, iannattone2020artificial]. These arteries branch out from the aorta and envelop the heart, forming an extensive network of vessels essential for cardiac function. Diseases affecting the coronary arteries can lead to myocardial ischemia, a condition characterized by reduced blood flow to the heart muscle [smirnova2023spontaneous, krittanawong2021machine, lionakis2022spontaneous]. Chronic ischemia or its associated complications can progressively impair the heart’s pumping efficiency, posing significant health risks [d2023deep]. Early detection and timely medical intervention are crucial for conditions related to coronary arteries, as they can be life-threatening if left unaddressed [lin2023coronary]. In the subsequent discussion, we delve into the realm of AI-driven analysis of diseases associated with coronary arteries. This exploration includes a comprehensive overview of the latest AI methodologies applied in the detection, diagnosis, and management of coronary artery diseases.
3.4.2.2 Coronary Artery Diseases (CAD)
CAD, the most prevalent form of heart disease, arises when coronary arteries are narrowed or blocked due to fatty plaque buildup [wahab2023developing]. This condition can significantly reduce or completely obstruct blood flow to the heart, leading to angina, manifested as chest pain or discomfort. CAD may also precipitate a heart attack, or myocardial infarction, where a part of the heart muscle sustains damage or death due to insufficient blood supply [zhang2022artificial]. Advanced stages of CAD often necessitate invasive treatments like Percutaneous Coronary Intervention (PCI) or coronary angioplasty, which involve widening narrowed arteries using a balloon device and potentially placing a stent to keep the artery open. In more severe cases with multiple blockages, Coronary Artery Bypass Grafting (CABG) surgery might be performed, creating alternative pathways for blood flow using vessels from other parts of the body [gautam2022current]. Timely and effective management of CAD is crucial. Below, we discuss significant research challenges associated with CAD. Over the years, AI techniques have become extensively utilized in the field of Coronary Artery Disease detection [alizadehsani2021coronary]. For example, Coronary Artery Stenosis is a predominant structural concern in coronary arteries. The detection, characterization, and monitoring of stenosis severity and plaque development are vital, as these conditions can lead to serious events like heart attacks [zreik2018recurrent]. Medical imaging is central to these tasks [mezquita2023clinical, li2001magnetic, van1995magnetic, pennell1996assessment, pennell1993magnetic, dhawan2004role, botnar2000noninvasive, van1999magnetic]. Notably, interpretations of coronary computed tomography angiography (CTA) by less experienced clinicians often overestimate the severity of stenosis compared to expert analyses [lin2022deep]. AI solutions present a promising alternative.